passing through (6, -3), and perpendicular to the line whose equation is x - 3y - 2 = 0 the equation of the line in point slope form is
Answer by mananth(16946)

-3 y = -1 x + 2
Divide by -3
y = 1/ 3 x + - 2/3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
Because m1*m2 =-1
The slope of the required line will be -3

m= -3 ,point ( 6 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = -18 + b
b= 15
m= -3
The required equation is y = -3 x + 15
m.ananth@hotmail.ca